OpenStudy (anonymous):

Solve the equation. If necessary, round your answer to the nearest tenth. x2 = 196 A. 14 B. –14, 14 C. 98 D. –98, 98

OpenStudy (zale101):

What number from these choices equals 196 when you square it?

OpenStudy (anonymous):

@Zale101 What

OpenStudy (zale101):

what part u didn't get exactly?

OpenStudy (anonymous):

The whole thing

OpenStudy (zale101):

the choices are A. 14 B.–14, 14 C.98 D. 98 14^2=? –14^2=?, 14^2=?. the result of -14 and 14 are the same. Because you are squaring means times it twice. Multiplying two negatives equals positive, multiplying two positives equal positive as well. -98^2=? 98^2=? both -98 and 98 will have the same result, and i already mentioned why. all you have to find is what's the square of 14 and 98. If one of them is 196 then it's the answer. The power of 2 means !4 X 14 = ? 98 X 98 = ? tell me what you get when you solve these, thank you

OpenStudy (zale101):

all u need is multiplying :)